2x=3y+11
7x-4y=6 Solve system with addition method

Answer by nabla(475)

First, multiply the first equation by -(7/2)
This gives:
-7x=(-21/2)y-77/2
7x=4y+6
Now we can add them.
0=(-13/2)y-65/2
65/2=(-13/2)y
-65/13=y
-5=y
Now for x:
7(x)+20=6
7x=-14
x=-2

check:
-4 =? -4
-14+20=6